how do I add and simplify 9u/(u^2-9)+u/(u-3)

Answer by jorel1380(3719)

9u/(u^2-9)+u/(u-3)=9u/(u+3)(u-3)+u(u+3)/(u+3)(u-3)=u2+12u/(u2-9)..